what is the pH of 1.003 M pyridine (Py and PyH+)?
Answer #1

V kol bar] - low 9.000 Port = -log Sko Lon] Pr - log St. LB7 = 14 H = 14+ log En Lolbo= 10003M K (199-1-7810-4 pH = 14.0 + log Skobo lx = 14.0 + lop 700 amx 1.7810-9 Pr = 14.0 + rog 117.05x1050 PH= 14.0+ log (